OpenStudy (anonymous):

can you explain in detail how to do the following question?: find the slope of a line parallel to the graph of each equation: 1. y=1/2x+2.3

OpenStudy (anonymous):

the slope is the same because parallel lines never intersect it the slope was different, they would intersect at one point

OpenStudy (anonymous):

Try the inverse

OpenStudy (anonymous):

To find the inverse, take y and x and interchange them and then solve for y. So the equation becomes x = 1/2y+2.3 Solve for y. x-2.3 = 1/2y x - 2.3*(2) = y y = 2(x -2.3) y = 2x - 4.6

OpenStudy (anonymous):

I read your question wrong! You need to find the y intercept and then make that your new equation's x intercept.

OpenStudy (anonymous):

um... please elaborate upon that...

OpenStudy (anonymous):

My answers aren't posting for some reason. You have this equation in the slope-intercept form already. The slope is 1/2. The slope of every parallel line is 1/2.